# High-Throughput and Scalable Secure Inference Protocols for Deep Learning with Packed Secret Sharing

### v1

@@ III-A Protocol Design
FC is a key layer in the deep learning §II-C, which consists of vector-matrix multiplications. In light of the earlier analysis, vector-matrix multiplication $\boldsymbol{a}\cdot A$ can be calculated by blocked matrix algorithms, where $\boldsymbol{a}\in\mathbb{F}_{p}^{u},A\in\mathbb{F}_{p}^{u\times v}$. Specifically, we divide the vector $\boldsymbol{a}\in\mathbb{F}_{p}^{u}$ into $\lceil u/k\rceil$ blocks and divide each column of the matrix $A\in\mathbb{F}_{p}^{u\times v}$ into $\lceil u/k\rceil$ blocks. The size of each block is $k$. Then we pack each block into a PSS (e.g., $\llbracket\boldsymbol{a}\rrbracket\in\mathbb{F}_{p}^{\lceil u/k\rceil},\llbracket A\rrbracket\in\mathbb{F}_{p}^{\lceil u/k\rceil\times v}$) and combine it with the DN [4] protocol.
@@
In order to reduce communication cost, we first define vector-matrix multiplication-friendly random share tuple (VM-RandTuple), i.e., $(\llbracket\boldsymbol{r}\rrbracket_{2d}\in\mathbb{F}_{p}^{k},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}\in\mathbb{F}_{p})$, where $\boldsymbol{r}=\{\boldsymbol{r}_{i}\}_{i=0}^{k^{2}-1},\boldsymbol{r}^{\prime}=\{\boldsymbol{r}^{\prime}_{i}=\sum_{j=i\cdot k}^{(i+1)\cdot k-1}\boldsymbol{r}_{j}\}_{i=0}^{k-1}$. We can utilize a Vandermonde matrix for the batch generation of such random shares tuple in the offline phase. We formally define the functionality $\mathcal{F}_{VM-RandTuple}$ in Appendix -A and provide a secure protocol implementation in Algorithm 1. Based on the VM-RandTuple, a key optimization to the second step of DN protocol is having $P_{1}$ combine these partial results into the final output and reshare the final output to the parties. This optimization directly leads to a $k$-fold reduction in the amortized communication cost of the second step of DN protocol, compared with [5].
@@
We describe a toy example in Fig. 3. It embodies a core idea of our technology. ① and ② correspond to general vector-matrix multiplication; ① and ③ correspond to fixed-point vector-matrix multiplication. We defer the detailed the latter to §III-B. In Fig. 3, we set $k=2$, meaning that $2$ values are packed into a single packed secret share. Accordingly, the vector $\boldsymbol{a}$ is packed into 2 packed secret shares, i.e., $\llbracket\boldsymbol{a}\rrbracket\in\mathbb{F}_{p}^{2}$ and the matrix $A$ is packed into $2\times 2$ packed secret shares, i.e., $\llbracket A\rrbracket\in\mathbb{F}_{p}^{2\times 2}$. Then all parties $\{P_{i}\}_{i=1}^{n}$ can compute $\llbracket\boldsymbol{a}\cdot A\rrbracket_{2d}=\llbracket\boldsymbol{a}\rrbracket_{d}\cdot\llbracket A\rrbracket_{d}$ locally. Since $\llbracket\boldsymbol{a}\cdot A\rrbracket_{2d}$ has a degree of $2d$, a degree reduction is required. $P_{i}$ sends the $\llbracket\boldsymbol{a}\cdot A\rrbracket_{2d}+\llbracket\boldsymbol{r}\rrbracket_{2d}$ to $P_{1}$ and $P_{1}$ reconstructs the secret values at the default positions $\{\boldsymbol{s}_{i}\}_{i=0}^{k-1}$ and then calculates the sum of the results. Finally, $P_{1}$ packs the multiple results, i.e., $(\boldsymbol{c}_{1},\boldsymbol{c}_{2})$ into a single degree-$d$ PSS and shares to the other parties. We formally define the functionality $\mathcal{F}_{VecMatMult}$ in Appendix -A and provide a secure protocol implementation in Algorithm 2.
@@
To sum up, given a vector of length $u$ and a $u$-by-$v$ matrix, the online complexity is exactly 1 round with $(1+\frac{1}{k})uv$ field elements per party. The offline complexity is exactly 2 round with $(1+\frac{1}{k})\frac{2nv}{n+2k-1}$ field elements per party.
@@
    There are $n$ parties where $n=2d+1,1<k\leq d$ and the number of corrupted servers is t=dk+1t=d-k+1. The original vector $\boldsymbol{a}\in\mathbb{F}_{p}^{u}$ and matrix $A\in\mathbb{F}_{p}^{u\times v}$.
@@
    Theorem 1.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
    Theorem 2.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
@@ III-B Truncation
    In secure inference, decimal numbers are encoded using fixed-point representation universally. In this representation, a decimal value is represented as $\ell$-bit integers, where the first $\ell-\ell_{x}$ bits are the integer part and the last $\ell_{x}$ bits are the fractional part. Multiplying two fixed-point numbers yields $2\ell_{x}$ fractional bits and $\ell-2\ell_{x}$ integer bits; truncation is required to reduce the fractional part back to $\ell_{x}$ bits, ensuring precision and avoiding overflow. [5] proposes a novel truncation protocol in the Mersenne prime field that reduces the gap between shares and the secret to just 1 bit. However, the new protocol is not well-suited for our work, as it incurs significant communication overhead. Specifically, the online communication overhead of the truncation protocol is the same as DN multiplication, but applying it in our work requires additional open operations and degree reduction. Therefore, we propose our fixed-point vector-matrix multiplication protocol that is a natural extension of ABY3 [9] in Algorithm 3. In Fig. 3, ① and ③ correspond to fixed-point vector-matrix multiplication.
@@
    Algorithm 3 Vector-Matrix Mult with Trunc $\Pi_{VecMatMult}^{Trunc}$
    We can see clearly that the key is how to generate the random shares $\llbracket\boldsymbol{r}\rrbracket_{2d},\llbracket\boldsymbol{r}^{\prime}\rrbracket_{d}$. The vectors $\boldsymbol{r}^{\prime}$ and $\boldsymbol{r}$ satisfy the following relations: $\boldsymbol{r}^{\prime}=\{\boldsymbol{r}^{\prime}_{i}\}^{k-1}_{i=0},\boldsymbol{r}=\{\boldsymbol{r}_{i}\}^{k^{2}-1}_{i=0},\boldsymbol{r}^{\prime}_{i}=(\sum_{i=j\cdot k}^{(j+1)\cdot k-1}\boldsymbol{r}_{i})/2^{\ell_{x}},j\in[k]$. We formally define the functionality $\mathcal{F}_{TruncTriple}$ in Appendix -A and provide a secure protocol implementation in Algorithm 4. Specifically, $\llbracket\boldsymbol{r}^{\prime}\rrbracket_{d}$ and $k$ random packed secret shares $\llbracket\boldsymbol{w}^{i}\rrbracket_{d},i\in[k]$ are genereted, where the length of vector $\boldsymbol{w}^{i}$ is $k$. Then there is the following relation: $\boldsymbol{r}_{j}=\boldsymbol{w}_{\lfloor j/k\rfloor}^{j\pmod{k}},j\in[k^{2}]$. Fig. 4 depicts a sample example. It should be noted that $\llbracket\boldsymbol{w}^{i}\rrbracket_{d}$ needs to be recombined. The recombination process consists of two steps: i) All secret values in vector $\boldsymbol{w}^{i}$ are converted into Shamir secret shares $\boldsymbol{w}^{i}_{j}|_{i}]_{d}$. Inspired by [10], we obtain Theorem 4. We treat a single packed secret share as $k$ distinct Shamir secret shares and apply Theorem 4 to achieve local conversion. We formalize this process in Algorithm 5; ii) These transformed values are recombined through linear combination with vector $\llbracket E_{i}\rrbracket_{k-1}$. Although Algorithm 5 performs the conversion of secret shares locally, it results in an increased degree of $2d$. Therefore, a degree reduction is required before recombination by functionality $\mathcal{F}_{DgreeTrans}$. We provide secure protocol implementation for $\mathcal{F}_{RandomBits},\mathcal{F}_{Random},\mathcal{F}_{DegreeTrans}$ in Appendix -B.
    Theorem 3.
    $\Pi_{VecMatMult}^{Trunc}$ securely realizes $\mathcal{F}_{VecMatMult}^{Trunc}$ in the $\mathcal{F}_{TruncTriple}$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
@@
    The proof of this theorem is given in Appendix -C. ∎
    Theorem 4.
@@
    The proof of this theorem is given in Appendix -C. ∎
@@
      Get $\boldsymbol{w}^{i}_{j}|_{s_{i}}]_{2d},i,j\in[k]$ by ShConvert($\llbracket\boldsymbol{w}^{i}\rrbracket_{d}$) (Algorithm 5) locally and get $\boldsymbol{w}^{i}_{j}|_{s_{i}}]_{2d-k+1}\leftarrow\mathcal{F}_{DegreeTrans}([\boldsymbol{w}^{i}_{j}|_{s_{i}}]_{2d})$.
@@
    Theorem 5.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
@@
    Algorithm 5 ShConvert
    To sum up, given a vector of length $u$ and a $u$-by-$v$ matrix, the online complexity is the same as protocol $\Pi_{VecMatMult}$. The offline complexity of this protocol depends on the generation of the random bits, random values and degree transformation, that is 5 rounds and $(n-1)(\frac{4\ell+k\ell+2k^{2}}{n-t}+\frac{4\ell+2k^{2}}{n}+k\ell+\ell)$ field elements per party for a truncation triple.
@@ IV-A Protocol Design
    FC is exactly a vector-matrix multiplication. Similarly, a Conv §II-C can also be expressed as a matrix multiplication that can be achieved through multiple vector-matrix multiplications. In light of the above analysis, when zero padding needs to be applied to convolutional input, the combination process incurs significant communication overhead.
    For avoiding significant communication cost, we present our efficient convolution computation protocol by leveraging the inherent parallelizability of convolution computation. To be specific, the input tensor or image is required to interact with every filter as shown in Fig. 2. Hence, the convolution process inherently involves parallel computation. In light of this, we propose a filter packing approach, i.e., we pack $k$ values that are selected respectively from the same position of $k$ filters into a single packed secret share. Correspondingly, we also pack $k$ copies of the input into a single packed secret share. Fig. 5 depicts a sample example. Then we can complete convolution computation by using a parallel matrix multiplication protocol. Further, when zero padding is applied to the input of convolutional operations, it can be performed locally without communication since we can directly pack $k$ zero values into a single packed secret share.
    We directly present the protocol for parallel fixed-point matrix multiplication in Algorithm 6 and formally define the functionality $\mathcal{F}_{PMatMult}^{Trunc}$ in Appendix -A. In offline Step 2, it holds $R_{\alpha,\beta}=\sum_{i=0}^{\ell-1}2^{i}\cdot R^{\prime\prime i}_{\alpha,\beta}\cdot R^{\prime\prime i}_{\alpha,\beta},$ since $R^{\prime\prime i}_{\alpha,\beta}\in\{0,1\}^{k}.$ In addition, $\llbracket\boldsymbol{0}\rrbracket_{d}$ can be generated by PRG without communication.
    To sum up, given a $u$-by-$v$ matrix and a $v$-by-$m$ matrix, the online complexity is 1 round with $\frac{2um}{k}$ field elements per party. The offline complexity is 4 rounds and $\frac{um\ell}{k}(\frac{3n}{n+2k-1}+2+\frac{n+2k}{2n})$ field elements per party.
@@
    Algorithm 6 Parallel Matrix Mult with Trunc $\Pi_{PMatMult}^{Trunc}$
    Theorem 6.
    $\Pi_{PMatMult}^{Trunc}$ securely realizes $\mathcal{F}_{PMatMult}^{Trunc}$ in the $\mathcal{F}_{RandomBits}$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎
@@ IV-B Inter-layer Connections
    When the next layer still performs a Conv operation, a pack transformation is required for the result of the convolution computation since we need to pack $k$ copies of the input into a single packed secret share. To be specific, the pack transformation is that given a PSS $\llbracket\boldsymbol{x}\rrbracket_{d}$, it outputs $\llbracket\boldsymbol{x}^{i}\rrbracket_{d}$, where $\boldsymbol{x}$ and $\boldsymbol{x}^{i}$ are two vectors of length $k,\boldsymbol{x}^{i}=(\boldsymbol{x}_{i},\cdots,\boldsymbol{x}_{i}),i\in[k]$. Fig. 5 presents an example. We formally define the functionality $\mathcal{F}_{PackTrans}$ in Appendix -A and provide a secure protocol implementation in Algorithm 7. In the offline, we generate $\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d},\llbracket\boldsymbol{r}\rrbracket_{d},$ where $\llbracket\boldsymbol{r}\rrbracket_{d}=(\boldsymbol{r}_{0},\cdots,\boldsymbol{r}_{k-1}),\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d}=(\boldsymbol{r}_{i},\cdots,\boldsymbol{r}_{i}),i\in[k].$ In the online, parties compute $\llbracket\boldsymbol{z}\rrbracket_{d}=\llbracket\boldsymbol{x}\rrbracket_{d}+\llbracket\boldsymbol{r}\rrbracket_{d}$ and reveal the vector $\boldsymbol{z}$ to $P_{1}$. Then, $P_{1}$ sets $\boldsymbol{z}^{i}=(\boldsymbol{z}_{i},\cdots,\boldsymbol{z}_{i})$ and distributes $\llbracket\boldsymbol{z}^{i}\rrbracket_{d}$. Finally, parties compute $\llbracket\boldsymbol{x}^{i}\rrbracket_{d}=\llbracket\boldsymbol{z}^{i}\rrbracket_{d}-\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d}$.
    To sum up, the online complexity is 1 round and $(n-1)(\frac{1}{2}+k)$ field elements per party. The offline complexity is 2 rounds and $(n-1)(\frac{2+k}{n-t}+\frac{2}{n})$ field elements per party.
@@
    Theorem 7.
    Proof.
    The proof of this theorem is given in Appendix -C. ∎

### v2

@@ III-A Protocol Design
FC is a key layer in the deep learning §II-C, which consists of vector-matrix multiplications. In light of the earlier analysis, vector-matrix multiplication $\boldsymbol{a}\cdot A$ can be calculated by blocked matrix strategy, where $\boldsymbol{a}\in\mathbb{F}_{p}^{u},A\in\mathbb{F}_{p}^{u\times v}$. Specifically, we divide the vector $\boldsymbol{a}\in\mathbb{F}_{p}^{u}$ into $\lceil u/k\rceil$ blocks and divide each column of the matrix $A\in\mathbb{F}_{p}^{u\times v}$ into $\lceil u/k\rceil$ blocks. The size of each block is $k$. Then we pack each block into a PSS (e.g., $\llbracket\boldsymbol{a}\rrbracket\in\mathbb{F}_{p}^{\lceil u/k\rceil},\llbracket A\rrbracket\in\mathbb{F}_{p}^{\lceil u/k\rceil\times v}$) and combine it with the DN [28] protocol.
@@
In order to reduce communication cost, we first define vector-matrix multiplication-friendly random share tuple (VM-RandTuple), i.e., $(\llbracket\boldsymbol{r}\rrbracket_{2d}\in\mathbb{F}_{p}^{k},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d}\in\mathbb{F}_{p})$, where $\boldsymbol{r}=\{\boldsymbol{r}_{i}\}_{i=0}^{k^{2}-1},\boldsymbol{r}^{\prime}=\{\boldsymbol{r}^{\prime}_{i}=\sum_{j=i\cdot k}^{(i+1)\cdot k-1}\boldsymbol{r}_{j}\}_{i=0}^{k-1}$. We can utilize a Vandermonde matrix for the batch generation of such random share tuples in the offline phase. We formally define the functionality $\mathcal{F}_{VM-RandTuple}$ in Appendix -C and provide a secure protocol implementation in Algorithm 1. Based on the VM-RandTuple, a key optimization to the second step of DN protocol is having $P_{1}$ combine these partial results into the final output and reshare the final output to the parties. This optimization directly leads to a $k$-fold reduction in the amortized communication cost of the second step of DN protocol, compared with LXY24.
@@
We describe a toy example in Fig. 3. It embodies a core idea of our technology. ① and ② correspond to general vector-matrix multiplication; ① and ③ correspond to fixed-point vector-matrix multiplication. We defer the details of the latter to §III-B. In Fig. 3, we set $k=2$, meaning that $2$ values are packed into a single packed secret share. Accordingly, the vector $\boldsymbol{a}$ is packed into 2 packed secret shares, i.e., $\llbracket\boldsymbol{a}\rrbracket\in\mathbb{F}_{p}^{2}$ and the matrix $A$ is packed into $2\times 2$ packed secret shares, i.e., $\llbracket A\rrbracket\in\mathbb{F}_{p}^{2\times 2}$. Then all parties $\{P_{i}\}_{i=1}^{n}$ can compute $\llbracket\boldsymbol{a}\cdot A\rrbracket_{2d}=\llbracket\boldsymbol{a}\rrbracket_{d}\cdot\llbracket A\rrbracket_{d}$ locally. Since $\llbracket\boldsymbol{a}\cdot A\rrbracket_{2d}$ has a degree of $2d$, a degree reduction is required. $P_{i}$ sends the $\llbracket\boldsymbol{a}\cdot A\rrbracket_{2d}+\llbracket\boldsymbol{r}\rrbracket_{2d}$ to $P_{1}$ and $P_{1}$ reconstructs the secret values at the default positions $\{\boldsymbol{s}_{i}\}_{i=0}^{k-1}$ and then calculates the sum of the results. Finally, $P_{1}$ packs the multiple results, i.e., $(\boldsymbol{c}_{1},\boldsymbol{c}_{2})$ into a single degree-$d$ PSS and shares to the other parties. We formally define the functionality $\mathcal{F}_{VecMatMult}$ in Appendix -C and provide a secure protocol implementation in Algorithm 2.
@@
To sum up, given a vector of length $u$ and a $u$-by-$v$ matrix, the online complexity is exactly 1 round with $(1+\frac{1}{k})uv$ field elements amortized per party. The offline complexity is exactly 2 round with $(1+\frac{1}{k})\frac{2nv}{n+2k-1}$ field elements amortized per party.
@@
    There are $n$ parties where $n=2d+1,1<k\leq d$ and the number of corrupted parties is t=dk+1t=d-k+1. The original vector $\boldsymbol{a}\in\mathbb{F}_{p}^{u}$ and matrix $A\in\mathbb{F}_{p}^{u\times v}$.
@@ III-B Truncation
    In MPC, decimal numbers are universally encoded using fixed-point representation. In this representation, a decimal value is represented as $\ell$-bit integers, where the first $\ell-\ell_{x}$ bits are the integer part and the last $\ell_{x}$ bits are the fractional part. Multiplying two fixed-point numbers yields $2\ell_{x}$ fractional bits and $\ell-2\ell_{x}$ integer bits; truncation is required to reduce the fractional part back to $\ell_{x}$ bits, ensuring precision and avoiding overflow. LXY24 proposes a novel truncation protocol in the Mersenne prime field that reduces the share-to-secret gap to just 1 bit, enabling the use of a 31-bit Mersenne prime. However, this protocol is not well-suited to our setting. First, integrating it into our work would require additional open operations and degree reduction. Second, for deeper neural networks, it necessitates either a 61-bit Mersenne prime or a reduction in precision to avoid overflow. Therefore, we propose a fixed-point vector-matrix multiplication protocol in the 61-bit Mersenne prime field, which serves as a natural extension of ABY3 [15], striking a balance between communication efficiency and precision in Algorithm 3. In Fig. 3, ① and ③ correspond to fixed-point vector-matrix multiplication.
@@
    Algorithm 3 Fixed Vector-Matrix Mult $\Pi_{VecMatMult}^{Fixed}$
    We can see clearly that the key is how to generate the random shares $\llbracket\boldsymbol{r}\rrbracket_{2d},\llbracket\boldsymbol{r}^{\prime}\rrbracket_{d}$. The vectors $\boldsymbol{r}^{\prime}$ and $\boldsymbol{r}$ satisfy the following relations: $\boldsymbol{r}^{\prime}=\{\boldsymbol{r}^{\prime}_{i}\}^{k-1}_{i=0},\boldsymbol{r}=\{\boldsymbol{r}_{i}\}^{k^{2}-1}_{i=0},\boldsymbol{r}^{\prime}_{i}=(\sum_{i=j\cdot k}^{(j+1)\cdot k-1}\boldsymbol{r}_{i})/2^{\ell_{x}},j\in[k]$. We formally define the functionality $\mathcal{F}_{TruncTriple}$ in Appendix -C and provide a secure protocol implementation in Algorithm 4. Specifically, $\llbracket\boldsymbol{r}^{\prime}\rrbracket_{d}$ and $k$ random packed secret shares $\llbracket\boldsymbol{w}^{i}\rrbracket_{d},i\in[k]$ are genereted, where the length of vector $\boldsymbol{w}^{i}$ is $k$. Then there is the following relation: $\boldsymbol{r}_{j}=\boldsymbol{w}_{\lfloor j/k\rfloor}^{j\pmod{k}},j\in[k^{2}]$. Fig. 4 depicts a sample example. It should be noted that $\llbracket\boldsymbol{w}^{i}\rrbracket_{d}$ needs to be recombined. The recombination process consists of two steps: i) All secret values in vector $\boldsymbol{w}^{i}$ are converted into Shamir secret shares $\boldsymbol{w}^{i}_{j}|_{i}]_{d}$. Inspired by [30], we obtain Theorem 2. We treat a single packed secret share as $k$ distinct Shamir secret shares and apply Theorem 2 to achieve local conversion. We formalize this process in Algorithm 5; ii) These transformed values are recombined through linear combination with vector $\llbracket E_{i}\rrbracket_{k-1}$. Although Algorithm 5 performs the conversion of secret shares locally, it results in an increased degree of $2d$. Therefore, a degree reduction is required before recombination by functionality $\mathcal{F}_{DgreeTrans}$. We provide secure protocol implementation for $\mathcal{F}_{RandomBits},\mathcal{F}_{Random},\mathcal{F}_{DegreeTrans}$ in Appendix -B.
    Proposition 1.
    If $xy\leq 2^{k}$ for some $k<\ell$, then with probability at least $1-2^{k}/(2^{\ell}-1)$ it holds that $z=\lfloor xy/2^{\ell_{x}}\rfloor+v$ for some $v\in\{0,1\}$.
@@
    The proof of this proposition is given in Appendix -D. ∎
    Proposition 2.
@@
    The proof of this theorem is given in Appendix -D. ∎
@@
      Get $\boldsymbol{w}^{i}_{j}|_{s_{i}}]_{2d},i,j\in[k]$ by $\Pi_{ShConvert}(\llbracket\boldsymbol{w}^{i}\rrbracket_{d})$(Algorithm 5) locally and get $\boldsymbol{w}^{i}_{j}|_{s_{i}}]_{2d-k+1}\leftarrow\mathcal{F}_{DegreeTrans}([\boldsymbol{w}^{i}_{j}|_{s_{i}}]_{2d})$.
@@
    Algorithm 5 $\Pi_{ShConvert}$
    To sum up, given a vector of length $u$ and a $u$-by-$v$ matrix, the online complexity is the same as protocol $\Pi_{VecMatMult}$. The offline complexity of this protocol depends on the generation of the random bits, random values and degree transformation, that is 5 rounds and $(n-1)(\frac{4\ell+k\ell+2k^{2}}{n-t}+\frac{4\ell+2k^{2}}{n}+k\ell+\ell)$ field elements amortized per party for a truncation triple.
@@ IV-A Protocol Design
    FC is exactly a vector-matrix multiplication. Similarly, a Conv §II-C can also be expressed as a matrix multiplication that can be achieved through multiple vector-matrix multiplications. In light of the above analysis, when zero padding needs to be applied to the convolutional input, the combination process incurs significant communication overhead.
    To avoid significant communication cost, we present our efficient convolution computation protocol by leveraging the inherent parallelizability of convolution computation. To be specific, the input is convolved with each filter in either a single-input multiple-output or multiple-input multiple-output convolution process as shown in Fig. 2. In light of this, we propose a filter packing approach, i.e., we pack $k$ values that are selected respectively from the same position of $k$ filters into a single packed secret share. Correspondingly, we also pack $k$ copies of the input into a single packed secret share. Fig. 5 depicts a sample example. Then we can complete convolution computation by using a parallel matrix multiplication protocol. Further, when zero padding is applied to the input of convolutional operations, it can be performed locally without communication since we can directly pack $k$ zero values into a single packed secret share.
    We directly present the protocol for parallel fixed-point matrix multiplication in Algorithm 6 and formally define the functionality $\mathcal{F}_{PMatMult}^{Fixed}$ in Appendix -C. In offline Step 2, it holds $R_{\alpha,\beta}=\sum_{i=0}^{\ell-1}2^{i}\cdot R^{\prime\prime i}_{\alpha,\beta}\cdot R^{\prime\prime i}_{\alpha,\beta},$ since $R^{\prime\prime i}_{\alpha,\beta}\in\{0,1\}^{k}.$ In addition, $\llbracket\boldsymbol{0}\rrbracket_{d}$ can be generated by PRG without communication.
    To sum up, given a $u$-by-$v$ matrix and a $v$-by-$m$ matrix, the online complexity is 1 round with $\frac{2um}{k}$ field elements amortized per party. The offline complexity is 4 rounds and $\frac{um\ell}{k}(\frac{3n}{n+2k-1}+2+\frac{n+2k}{2n})$ field elements amortized per party.
@@
    Algorithm 6 Fixed Parallel Matrix Mult $\Pi_{PMatMult}^{Fixed}$
@@ IV-B Inter-layer Connections
    When the next layer still performs a Conv operation, a pack transformation is required for the result of the convolution computation since we need to pack $k$ copies of the input into a single packed secret share. To be specific, the pack transformation is that given a PSS $\llbracket\boldsymbol{x}\rrbracket_{d}$, it outputs $\llbracket\boldsymbol{x}^{i}\rrbracket_{d}$, where $\boldsymbol{x}$ and $\boldsymbol{x}^{i}$ are two vectors of length $k,\boldsymbol{x}^{i}=(\boldsymbol{x}_{i},\cdots,\boldsymbol{x}_{i}),i\in[k]$. Fig. 5 presents an example. We formally define the functionality $\mathcal{F}_{PackTrans}$ in Appendix -C and provide a secure protocol implementation in Algorithm 7. In the offline, we generate $\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d},\llbracket\boldsymbol{r}\rrbracket_{d},$ where $\llbracket\boldsymbol{r}\rrbracket_{d}=(\boldsymbol{r}_{0},\cdots,\boldsymbol{r}_{k-1}),\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d}=(\boldsymbol{r}_{i},\cdots,\boldsymbol{r}_{i}),i\in[k].$ In the online, parties compute $\llbracket\boldsymbol{z}\rrbracket_{d}=\llbracket\boldsymbol{x}\rrbracket_{d}+\llbracket\boldsymbol{r}\rrbracket_{d}$ and reveal the vector $\boldsymbol{z}$ to $P_{1}$. Then, $P_{1}$ sets $\boldsymbol{z}^{i}=(\boldsymbol{z}_{i},\cdots,\boldsymbol{z}_{i})$ and distributes $\llbracket\boldsymbol{z}^{i}\rrbracket_{d}$. Finally, parties compute $\llbracket\boldsymbol{x}^{i}\rrbracket_{d}=\llbracket\boldsymbol{z}^{i}\rrbracket_{d}-\llbracket\boldsymbol{r}^{\prime i}\rrbracket_{d}$.
    To sum up, the online complexity is 1 round and $(n-1)(\frac{1}{2}+k)$ field elements amortized per party. The offline complexity is 2 rounds and $(n-1)(\frac{2+k}{n-t}+\frac{2}{n})$ field elements amortized per party.